OpenStudy (anonymous):

7z - 1/2 = 3z - 2 z=

OpenStudy (anonymous):

First get all of your z terms on one side and all of your numbers without z on the other, so add 1/2 to both sides and subtract 3z from both sides to get: 4z=-3/2 Then divide both sides by 4 and you get: z=-3/8

OpenStudy (anonymous):

I did a whole bunch of research online with this question, and they all had that answer, but the answer box wont let me put more than 2 numbers in it

OpenStudy (anonymous):

Z=\[\frac{ -3 }{8 }\] Simplify both sides of the equation. Subtract 3z from both sides. Add 1/2 to both sides. Divide both sides by 4.

OpenStudy (anonymous):

thank you both xx is there a way to write -3/8 as a decimal?